the line through (6, -4) and (-3, -2) is parallel to the line through (2, 1) and (0, y). find y.
Answer by mananth(16949)

x1 y1 x2 y2
6 -4 -3 -2

slope m = (y2-y1)/(x2-x1)
( -2 - -4 )/( -3 -
( 2 / -9 )
m= - 2/ 9

slope of the parallel line will be the same.
(2, 1) and (0, y). find y.
-2/9 = (y-1)/ ( 0-2)
-2/9 = y-1/-2
4/9 = y-1
y=4/9 + 1
y=13/9
Happy New Year
m.ananth@hotmail.ca